I need to find the LCM of the following Algebraic Expressions:
x^5+2x^4+x^3, 2x^3-2x, 5x-5
This is how I solved the problem. Am I correct?
x^5+2x^4+x^3
x^3(x^2+2x+1)
x^3(x+1)(x+1)
2x(x^2-1)
2x(x+1)(x-1)
5x-5
5(x-1)
x^3(x+1)(x+1)2x(x+1)(x-1)5(x-1)
LCM= 10x^4(x+1)^2(x-1)

Answer by Alan3354(69443)

I need to find the LCM of the following Algebraic Expressions:
x^5+2x^4+x^3, 2x^3-2x, 5x-5
This is how I solved the problem. Am I correct?
x^5+2x^4+x^3
x^3(x^2+2x+1)
x^3(x+1)(x+1)
2x(x^2-1)
2x(x+1)(x-1)
5x-5
5(x-1)
x^3(x+1)(x+1)2x(x+1)(x-1)5(x-1)
LCM= 10x^4(x+1)^2(x-1)
-----------
x^5+2x^4+x^3 = x^3(x+1)(x+1)
2x^3-2x = 2x(x+1)(x-1)
5x-5 = 5(x-1)
----------------
LCM = x^3(x+1)(x+1)*2*(x-1)*5
= 10x^3(x+1)^2(x-1)
--------------
I don't think you need x^4, just x^3